OpenStudy (anonymous):

Simplify. 4 × 32 + 3 ÷ 3 - (1.1 + 5.9) 30 41.8 40 20

OpenStudy (anonymous):

use order of operations

OpenStudy (anonymous):

weird, if i use order of operations i dont get any of the choices

OpenStudy (anonymous):

oh wait

OpenStudy (anonymous):

fml figured it out

OpenStudy (anonymous):

actually

OpenStudy (anonymous):

according to google calculator, the answer should be 122

OpenStudy (anonymous):

That's not very helpful. You should've showed the steps better, not the direct answer

OpenStudy (anonymous):

im not giving the answer seeing that the answer is none of the choices

OpenStudy (anonymous):

im just pointing out that theres something wrong

OpenStudy (anonymous):

Maybe try using the drawing tool to show step-by-step to show to the user

OpenStudy (anonymous):

go ahead you can im not. it will be a waste of time

OpenStudy (anonymous):

goodbye

OpenStudy (anonymous):

Then why are you answering?

OpenStudy (anonymous):

I'm just a watch account. Once school starts again I will be making a real to account to answer. Sorry!

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

@mathmate

OpenStudy (anonymous):

@SneliS

OpenStudy (anonymous):

I believe 122

OpenStudy (nerdgirl):

Alright, so the order of operations is as follows: PEMDAS. and i'm sure you know what all of that stands for, right?

OpenStudy (anonymous):

not 32 its 3 and the 2 above it

OpenStudy (anonymous):

@SneliS

OpenStudy (anonymous):

facepalm that changes everything

OpenStudy (anonymous):

it would be 3^2

OpenStudy (anonymous):

it that case it would be A

OpenStudy (mathivh):

\[(4 × 3^2) + (3 ÷ 3) - (1.1 + 5.9)\]\[=(36) + (1) - (7) = 30\]. Thus the answer is A.

OpenStudy (anonymous):

PEMDAS Parentheses first so 5.9+1.1 is 7 then exponents so 3^2 is 9 multiplication so 9 x 4 is 36 division 3 divided by 3 is 1 addition 36+1 is 37 subtraction 37-7 is 30 30 is your answer

OpenStudy (anonymous):

ye xd

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

Do you still need help? @jordanjamesbay

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

That is much better teamwork.

OpenStudy (anonymous):

Alright. Allow me to show you. :D

OpenStudy (anonymous):

Use the Order of Operations: \[4~×~3^2~+~3~÷~3~-~(1.1~+~5.9) \]\[4~×~3^2~+~3~÷~3~-~(7) \]\[36~+~3~÷~3~-~(7)\]\[36~+~1~-~(7)\]\[37~-~(7)\]\[30\] Do you understand now? :)

OpenStudy (anonymous):

@jordanjamesbay Did you see my answer. :D

OpenStudy (anonymous):

ya